40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10035     Entry time: Fri Jun 13 09:20:37 2014
Author: Steve 
Type: Update 
Category: SUS 
Subject: restored damping at PRM and ETMY 

ETMX medm screen values are blank.

ELOG V3.1.3-